Finding Area by the Limit Definition In Exercises 47–56, use the limit process to find the area of the region bounded by the only the ones highlighted and please show all steps.
Answer #1

47 9 Area (A) aza lim din is y=0) stories n200 Ar= 6-9 y=f(1) =-44775 [01] s finida Limit definition of staid Ĉ famigax. = 1-